OpenStudy (anonymous):

The male to female ratio in taiwan is 52 to 48 write this ratio as a fraction in reduced form

OpenStudy (cwrw238):

52 = 2 * 2* 13 48 = 2*2*2*2*3 cancel ut the factors which are common

OpenStudy (anonymous):

\[\frac{52}{48}=\frac{13}{12}\]

OpenStudy (anonymous):

thanks